OpenStudy (anonymous):

Solving Systems of Equations by Substitution. y= -3x + 5 5x - 4y = -3

OpenStudy (p0sitr0n):

the goal of substitution is to replace one variable`s value by another so, 5x -4(-3x +5) +3 =0 5x +12x-20 + 3 = 0 17x -17 =0 17x = 17 x =1

OpenStudy (p0sitr0n):

y =2